# Gravitational Lensing of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter

### v1

@@ 3.3 Time Delay of Light
In the gravitational field, the elapse of time (measured by an observer) during the light propagation can be strongly affects by gravitation, and this is the time delay effect of gravitational field. To calculate the time delay of light in the gravitational lensing observation, it is necessary to computation the variation of time coordinate $t$ (measured by an distant observer) in the photon orbit. For the massless photon traveling along null geodesics (with $\epsilon=0$), one can derive the time derivative of radial coordinate from the reduced geodesic equation in (9)
@@
The plus and minus sign $\pm$ can be determined similar to the calculations of gravitational deflection angle. When the photon beams emitted from the source position $r_{\text{S}}$, the radial coordinate $r$ decreases as time passes, until the photon beams reach the closest distance $r=r_{0}$ to central supermassive black hole. After passing the turning point $r=r_{0}$, the radial coordinate starts to increase as time advances.
@@
In the gravitational lensing observation, when the light source and observer are located at $r=r_{\text{S}}$ and $r=r_{\text{O}}$, the time delay of light during the propagation process can be expressed as
@@
with $T_{0}=\sqrt{r_{\text{S}}^{2}-r_{0}^{2}}+\sqrt{r_{\text{O}}^{2}-r_{0}^{2}}$ to be the time period during the light propagation process without the gravitational field. In the expression (18), it is obvious that the time delay $\Delta T$ increases monotonically as the coordinates of light source and observer $r_{\text{S}}$, $r_{\text{O}}$ increase. In the integration process, the turning point $r=r_{0}$ is solved through equation (14).
Furthermore, it should be addressed that the time delay of light calculated using expression (18) is the time delay measured by the conventional static time coordinate $t$ in spacetime metric expression (7). For an asymptotically flat spacetime, the static time coordinate corresponds to the time measured by a very distant observer. On the other hand, if one want to calculate the time delay measured by the proper time of an finite distance static observer, a redshift factor should be taken into account
$\Delta T_{\text{observer's proper time unit}}=\frac{\Delta T}{\text{redshift factor}}=\frac{\Delta T}{\sqrt{f(r_{\text{O}})}}$ (19)
@@
3.4 Massive Object Orbits and Precession
Besides the gravitational deflection discussed in subsection 3.2 (which is relevant to a scattering orbit for massless photons), the massive object’s orbits are also greatly influenced by gravitational field, both for bound orbits and scattering orbits. Notably, massive object’s bound orbits may exhibit precession. In the gravitational lensing observations, the precession of bound orbits is frequently used to test the gravitational field around the supermassive black hole in galaxy center (such as the observation of Sgr*A in Milky Way Galaxy).
@@ 3.4 Massive Object Orbits and Precession
where we have used the new variable $u=1/r$ in the geodesic equation. Differentiating each side of this equation with respect to $\phi$, a second order differential equation for massive object’s orbit can be acquired
@@
An important feature of the massive object’s bound object is the precession of azimuthal angle $\phi$, which is illustrated in figure 2. Given a bound timelike geodesic orbit for a massive object, there are maximal and minimal position to central black hole in this orbit, denoting as $r_{\text{max}}$ and $r_{\text{min}}$. The precession of azimuthal angle $\phi$ in an orbit period can be calculated as
@@
The plus and minus sign $\pm$ can be determined as follows. When a massive object begin to move along timelike geodesics from the maximal distance position $r_{\text{max}}$, the radial coordinate $r$ decreases as azimuthal angle $\phi$ increases, until the massive object reaches the minimal distance position $r=r_{\text{min}}$ to central supermassive black hole. After passing the minimal distance position $r=r_{\text{min}}$, the massive object’s radial coordinate starts to increase as azimuthal angle increases.
@@ 4 Results and Discussions
This section presents the results on the gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM, including the gravitational deflection angle of light, time delay of light, the trajectory and precession angle of massive object’s bound orbits. These results are calculated numerically using expressions (15), (18) and (22). Furthermore, inspired by the rapid development of black hole shadow since the capture of black hole images by ETH, we also present a discussion on the black hole shadow for Euler-Heisenberg black hole surrounded by PFDM.
@@ 4.1 Gravitational Deflection Angle of Light
Figure 3: The gravitational deflection angle of light for Euler-Heisenberg black hole surrounded by PFDM. The gravitational deflection angles are calculated numerically from expression (15) for finite distance light source and observer. This figure highlights the influences of black hole electric charge, nonlinear electrodynamics effects and dark matter effects on gravitational deflection angles. (a) The upper left panel shows the influences from black hole electric charge, with the nonlinear electrodynamics parameter $a=0.5M^{2}$ and PFDM parameter $\lambda_{\text{DM}}=-0.2M$. (b) The upper right panel shows the influences from nonlinear electrodynamics effects / QED effects, with the black hole charge $Q=0.5M$ and PFDM parameter $\lambda_{\text{DM}}=-0.2M$. (c) The lower left panel highlights the influences form dark matter (for several PFDM parameter $\lambda_{\text{DM}}$ values) in the absence of electric charge and nonlinear electrodynamics effects ($Q=0$ and $a=0$). (d) The lower right panel highlights the influences form dark matter in the presence of nonlinear electrodynamics effects, with the black hole charge $Q=0.5M$ and nonlinear electrodynamics parameter $a=0.5M^{2}$. In all four panels of this figure, the horizontal axis labels the impact parameter in photon orbit, which is increased form $b=10M$ to $b=100M$. The position of light source and observer are located at a certain distance $r_{\text{O}}=r_{\text{S}}=200M$.
@@
The gravitational deflection angle of light for Euler-Heisenberg black hole surrounded by PFDM are given in figure 3. They are calculated numerically from expression (15) for light source and observer located at finite distance region to the central black hole. In figure 3, the upper left and upper right panels highlight the influences from black hole electric charge $Q$ and the nonlinear electrodynamics effects / QED effects (parameterized by $a$), while the lower left and lower right panels highlight the influences form dark matter parameterized by $\lambda_{\text{DM}}$. This figure shows the gravitational deflection in strong gravitational field cases, with the minimal impact parameter value is $b\sim 10M$, which is closer to the black hole critical shadow radius.
@@
The upper left panel of figure 3 suggests that, for a constant PFDM parameter and non-linear electrodynamics parameter, the gravitational deflection angles are slightly reduced when black hole’s electric charge become larger. But the effect from electric charge is small such that we must add a magnified subplot into upper left panel to separate the various curves corresponding different black hole charges. The upper right panel of figure 3 suggests that the variation of nonlinear electrodynamics parameter has no obvious influences on the gravitational deflection angles, even if a magnified subplot is employed. Comparing the upper and lower panels of figure 3, we can observe that both electric charge and nonlinear electrodynamics have minor impacts on gravitational deflection angle, while the most significant influences coming from the PFDM. From the lower panels of figure 3, it can be clearly manifested that the gravitational deflection angle grow rapidly as the increasing of PFDM parameter $\lambda_{\text{DM}}$. Notably, when PFDM parameter becomes negative, the gravitational deflection angle could grow to a very large value. This can be explained from the spacetime metric in expression (6) for Euler-Heisenberg black hole surrounded by PFDM. The dark matter effects contribute to the spacetime metric function $f(r)$ via the $-\lambda_{\text{DM}}/r$ term (if we naively omit the logarithmic term of PFDM for simplicity), which is similar to the black hole’s mass term $2M/r$ (both of them are inverse proportional to the radial coordinate). This indicates that the PFDM could play a role of “effective mass” in the spacetime metric (specifically specking, because of the presence of logarithmic term, this “effective mass” should be $m_{\text{eff}}(r)=\lambda_{\text{DM}}\ln\frac{r}{|\lambda|}$, rather than $m_{\text{eff}}(r)=\lambda_{\text{DM}}$). A negative PFDM parameter ($\lambda_{\text{DM}}<0$) contributes to a positive effective mass in the spacetime metric, which explains the rapid growth of the gravitational deflection angle for negative PFDM parameter values.
@@ 4.2 Time Delay of Light
Figure 4: The gravitational time delay of light for Euler-Heisenberg black hole surrounded by PFDM. The time delays of light (which are measured by the conventional static time coordinate $t$) are calculated numerically from expression (18) for finite distance light source and observer. This figure plots the time delay results measured by conventional static time coordinate $t$, highlighting the influences coming from various black hole parameters. (a) The upper left panel shows the influences from black hole electric charge on time delay, with the nonlinear electrodynamics parameter $a=0.5M^{2}$ and dark matter parameter $\lambda_{\text{DM}}=-0.2M$. (b) The upper right panel shows the influences from nonlinear electrodynamics effects / QED effects, with the black hole charge $Q=0.5M$ and PFDM parameter $\lambda_{\text{DM}}=-0.2M$. (c) The lower left panel highlights the influences form dark matter (for several PFDM parameter $\lambda_{\text{DM}}$ values) in the absence of black hole charge and nonlinear electrodynamics effects ($Q=0$ and $a=0$). (d) The lower right panel highlights the influences form dark matter in the presence of nonlinear electrodynamics effects, with the black hole charge $Q=0.5M$ and nonlinear electrodynamics parameter $a=0.5M^{2}$. In all four panels of this figure, the horizontal axis labels the position of light source varied from $r_{\text{S}}=100M$ to $r_{\text{S}}=1000M$, and the vertical axis is the time delay of light $\Delta T$ in unit of black hole mass. The position of observer is chosen to be $r_{\text{O}}=200M$, and the impact parameter in photon orbit satisfies $b=10M$.
@@
The time delays of light for Euler-Heisenberg black hole surrounded by PFDM are presented in figure 4. This figure plot the time delay results measured by the conventional static time coordinate $t$ in spacetime metric (7). They are obtained numerically from expression (18) for light source and observer located at a finite distance region. In this figure, we plot the time delay of light varied with respect to light source position $r_{\text{S}}$. A similar plot can be carried out for time delay varied with respect to observer position $r_{\text{O}}$, while the results exhibit very similar behaviors with those in figure 4. So the plot of time delay varied with respect to observer position is not shown in this work. In figure 4, the upper left and upper right panels highlight the influences from black hole electric charge $Q$ and the nonlinear electrodynamics effects (or QED effects) parameterized by $a$, while the lower left and lower right panels highlight the influences form dark matter parameterized by $\lambda_{\text{DM}}$. In all four panels, the time delays of light all grow with the increasing of light source radius $r_{\text{S}}$, indicating that the time delay of light $\Delta T$ calculated using expression (18) is a monotonically function of light source radius $r_{\text{S}}$ (or observer radius $r_{\text{O}}$).
@@
Comparing the time delay results in figure 4 and the gravitational deflection results in figure 3, it is easy to observe that the black hole parameters (including the black hole electric charge $Q$, nonlinear electrodynamics parameter $a$, PFDM parameter $\lambda_{\text{DM}}$) influence time delay of light in a similar trend as those for gravitational deflection angles. Firstly, Euler-Heisenberg black hole with a larger electric charge results in a reduction of time delay, as shown in the upper left panel of figure 4. Secondly, the nonlinear electrodynamics effects (parameterized by $a$) do not have obvious influences on time delay results, as indicated in the upper right panel of figure 4. Thirdly, the most significant impacts on time delay coming from the PFDM, which are much greater than those from electric charge and non-linear electrodynamics. From the lower panels of figure 4, the time delay of light changes rapidly as the PFDM parameter changes. Particularly, the time delay of light could grow to a very large value for a negative PFDM parameter ($\lambda_{\text{DM}}<0$), compared with those in the absence of dark matter ($\lambda_{\text{DM}}=0$). This can be explained by the “effective mass” role of PFDM in the spacetime metric, the same as we have elaborated in subsection 4.1 for gravitational deflection angle cases. A negative PFDM parameter contributes to a positive effective mass in the spacetime metric, which explains the rapid growth of the time delay for negative PFDM parameter values.
4.3 Massive Object Orbits and Precession
@@
Figure 5: The precession angles and trajectories of massive particle orbits moving around the Euler-Heisenberg black hole in the presence of PFDM. The precession angles $\Delta\phi_{\text{precession}}$ plotted in this figure are calculated numerically from expression (23) for a given conserved angular momentum value, and the trajectories of orbits are obtained by numerically solving the differential equation (22). The left panels shows the variation of precession angle as the total energy parameter $\varepsilon_{\text{energy}}=\frac{E^{2}-1}{2}$ changes, while the right panels illustrate the trajectories of particle orbits for a given total energy parameter (It should be noted that here we have substituted $1/2$ in the right hand side of equation (9) such that the energy parameter vanished at infinity, namely $\lim_{r\to\infty}\varepsilon_{\text{energy}}=0$). (a) The upper left panel plots the precession angle of orbits affects by black hole charge $Q$. (b) The upper right panel illustrates the trajectories of orbits for different black hole charge values when total energy equals $\varepsilon_{\text{energy}}=-0.03$ (c) The middle left panel shows the precession angle of orbits affects by dark mater (parameterized by $\lambda_{\text{DM}}$). (d) The middle right panel illustrates the trajectories of orbits for different PFDM parameter values when total energy is $\varepsilon_{\text{energy}}=-0.005$ (e) The lower left panel shows the precession angle of orbits influenced by nonlinear electrodynamics effects / QED effects (parameterized by $a$). (f) The lower right panel illustrates the trajectories of orbits for different nonlinear electrodynamics parameter values when total energy equals $\varepsilon_{\text{energy}}=-0.03$.
@@
The figure 5 presents results on massive object’s bound orbits around the Euler-Heisenberg black hole, including the precession angles in the left panels and the trajectories of orbits in the right panels. The trajectories of massive object’s orbits are obtained by numerically solving the ordinary differential equation (22), and the precession angle of orbits are calculated from expression (23). Correspondingly, the upper, middle and lower panels of figure 5 highlight the influences from black hole electric charge, dark matter (parameterized by PFDM parameter $\lambda_{\text{DM}}$), and nonlinear electrodynamics effects / QED effects (parameterized by nonlinear electrodynamics parameter $a$). In this figure, we present results on the trajectory and precession angle of massive object’s orbits for a selected angular momentum value $J=8M$. The precession angles correspond to other angular momentum values are similar to those for $J=8M$ given in figure 5.
@@
The results in figure 5 show that the black hole electric charge and PFDM can intensely influence the precession angle of massive object’s orbits in the vicinity of Euler-Heisenberg black hole. The upper panels of figure 5 suggests that a larger black hole charge reduces the precession angle of orbits. However, unlike the precession angle, the shape of massive object’s bound orbits are not violently influenced by black hole charge. In the upper right panel of figure 5, the shape and eccentricity of massive object’s orbits calculated for different black hole charge values do not exhibit large differences. On the other hand, the middle panels indicate that the PFDM have great impacts not only on the precession angle of orbits, but also on the shape of massive object’s bound orbits. The middle left panel shows that a negative PFDM parameter can significantly enhance the precession angle. The middle right panel shows that the shape, eccentricity and semimajor-axis length of massive object’s bound orbits (when PFDM parameter takes different values) exhibit notable differences. Moreover, in the middle left panel, the energy parameter $\varepsilon_{\text{energy}}=\frac{E^{2}-1}{2}$ could not cover the whole region $-0.03\leq\varepsilon_{\text{energy}}\leq 0$ when PFDM parameter $\lambda_{\text{DM}}>-0.4M$. This is due to the local minimum position of effective potential for massive objects. For any massive object’s bound orbits with precession, the total energy $\varepsilon_{\text{energy}}$ in this orbit must be greater than the local minimum of effective potential. The local minimum values of effective potential for massive objects moving around Euler-Heisenberg black hole in the presence of PFDM (for a selected angular momentum value $J=8.0M$ and several PFDM parameter values) are listed in table 1. Unlike the black hole electric charge and PFDM, the nonlinear electrodynamics effects / QED effects (parameterized by $a$) have very tiny impacts on the trajectories of massive object’s orbits and their procession angles, which can be directly observed from the lower panels of figure 5. This trend is similar to the cases of gravitational deflection and time delay, where the nonlinear electrodynamics effects have very tiny influences on these gravitational lensing observables.

### v2

@@ 3.3 Time Delay of Light
In the gravitational field, the elapse of time (measured by an observer) during the light propagation can be strongly affected by gravitation, and this is the time delay effect of gravitational field. To calculate the time delay of light in gravitational lensing observations, it is necessary to compute the variation of time coordinate $t$ (measured by an distant observer) in the photon orbit. For the massless photon traveling along null geodesics (with $\epsilon=0$), one can derive the time derivative of radial coordinate from the reduced geodesic equation in (10)
@@
The plus and minus sign $\pm$ can be determined similar to the calculation of gravitational deflection angle. When the photon beams are emitted from the source position $r_{\text{S}}$, the radial coordinate $r$ decreases as time passes, until the photon beams reach the closest distance $r=r_{0}$ to central supermassive black hole. After passing the turning point $r=r_{0}$, the radial coordinate starts to increase as time advances. So we have the relation
@@
In the gravitational lensing observation, when the light source and observer are located at $r=r_{\text{S}}$ and $r=r_{\text{O}}$, the time delay of light during the light propagation process can be expressed as
@@
with $T_{0}=\sqrt{r_{\text{S}}^{2}-r_{0}^{2}}+\sqrt{r_{\text{O}}^{2}-r_{0}^{2}}$ to be the time period during the light propagation process without the gravitational field. In the expression (19), it is obvious that the time delay $\Delta T$ increases monotonically as the coordinates of light source and observer $r_{\text{S}}$, $r_{\text{O}}$ increase. Additionally, in the integration, the turning point $r=r_{0}$ is solved through equation (15).
Furthermore, it should be addressed that the time delay of light calculated using expression (19) is the time delay measured by the conventional static time coordinate $t$ in spacetime metric expression (8). For an asymptotically flat spacetime, the static time coordinate corresponds to the time measured by a very distant observer (at the asymptotically flat region). On the other hand, if one wants to calculate the time delay measured by the proper time of an finite distance static observer, a redshift factor should be taken into account
$\Delta T_{\text{static observer's proper time unit}}=\frac{\Delta T}{\text{redshift factor}}=\frac{\Delta T}{\sqrt{f(r_{\text{O}})}}$ (20)
@@
3.4 Massive Object’s Orbits and Precession Angle
Besides the gravitational deflection angle and time delay of light discussed in subsections 3.2-3.3 (which is relevant to a scattering orbit for massless photons), the massive object’s orbits are also greatly influenced by gravitational field, both for bound orbits and scattering orbits. Notably, massive object’s bound orbits may exhibit precession. In the gravitational lensing observations, the precession of bound orbits is frequently used to test the gravitational field around the supermassive black hole in galaxy center (such as the observations relevant to Sgr*A in Milky Way Galaxy).
@@ 3.4 Massive Object’s Orbits and Precession Angle
where we have used the new variable $u=1/r$ in the geodesic equation. Differentiating each side of this equation with respect to $\phi$, a second order differential equation for massive object’s orbit can be obtained
@@
An important feature of the massive object’s bound object is the precession of azimuthal angle $\phi$, which is illustrated in figure 2. Given a bound timelike geodesic orbit for a massive object, there are maximal and minimal distance positions to central black hole in this orbit, denoting as $r_{\text{max}}$ and $r_{\text{min}}$. The precession of azimuthal angle $\phi$ in an orbit period can be calculated as
@@
The plus and minus sign $\pm$ can be determined as follows. When a massive object begins to move along timelike geodesics from the maximal distance position $r_{\text{max}}$, the radial coordinate $r$ decreases as azimuthal angle $\phi$ increases, until the massive object reaches the minimal distance position $r=r_{\text{min}}$ to central supermassive black hole. After passing the minimal distance position $r=r_{\text{min}}$, the massive object’s radial coordinate starts to increase as azimuthal angle increases. Thus we can get the relation
@@ 4 Results and Discussions
This section presents results on the gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM, including the gravitational deflection angle of light, time delay of light, the trajectory and precession angle of massive object’s bound orbits. These results are calculated numerically using expressions (16), (19) and (23). Furthermore, inspired by the rapid development of black hole shadow researches since the capture of black hole images by ETH, we also present a discussion on the black hole shadow for Euler-Heisenberg black hole surrounded by PFDM.
@@ 4.1 Gravitational Deflection Angle of Light
Figure 3: The gravitational deflection angle of light for Euler-Heisenberg black hole surrounded by PFDM. The gravitational deflection angles are calculated numerically from expression (16) for finite distance light source and observer. This figure highlights the influences of black hole electric charge, nonlinear electrodynamics effects and dark matter effects on gravitational deflection angles. (a) The upper left panel exhibits the influences from black hole electric charge $Q$, with the nonlinear electrodynamics parameter $a=0.5M^{2}$ and PFDM parameter $\lambda_{\text{DM}}=0.2M$. (b) The upper right panel shows the influences from nonlinear electrodynamics effects / QED effects, with the black hole charge $Q=0.5M$ and PFDM parameter $\lambda_{\text{DM}}=0.2M$. (c) The lower left panel highlights the influences form dark matter in the absence of electric charge and nonlinear electrodynamics effects (with $Q=0$ and $a=0$). (d) The lower right panel highlights the influences form dark matter in the presence of nonlinear electrodynamics effects, with the black hole charge $Q=0.5M$ and nonlinear electrodynamics parameter $a=0.5M^{2}$. In all four panels of this figure, the horizontal axis labels the impact parameter in photon orbit, which is increased form $b=10M$ to $b=100M$. The position of light source and observer are located at a certain distance $r_{\text{O}}=r_{\text{S}}=200M$.
@@
The gravitational deflection angle of light for Euler-Heisenberg black hole surrounded by PFDM are given in figure 3. They are calculated numerically from expression (16) for light source and observer located at finite distance region. In figure 3, the upper left and upper right panels highlight the influences from black hole electric charge $Q$ and the nonlinear electrodynamics effects / QED effects (parameterized by $a$), while the lower panels highlight the influences form dark matter parameterized by $\lambda_{\text{DM}}$. To shows the gravitational deflection in strong gravitational field cases, the minimal value of impact parameter is chosen as $b\sim 10M$, which indicates the photon orbit is very close to the black hole critical shadow radius.
@@
The upper left panel of figure 3 suggests that, for a constant PFDM parameter and nonlinear electrodynamics parameter, the gravitational deflection angles are slightly reduced when black hole’s electric charge become larger. But the effect coming from electric charge is small such that we must add a magnified subplot into upper left panel to separate the various curves corresponding to different black hole charges. The upper right panel of figure 3 suggests that the nonlinear electrodynamics parameter has no obvious influences on the gravitational deflection angles, even if a magnified subplot is employed. Comparing the upper and lower panels of figure 3, we can observe that both electric charge and nonlinear electrodynamics have minor impacts on gravitational deflection angle, while the most significant influences coming from the PFDM. From the lower panels of figure 3, it can be clearly manifested that the variation of PFDM parameter could dramatically change the gravitational deflection angle of light for charged Euler-Heisenberg black hole. Notably, the gravitational deflection angle reduce rapidly as the increasing of PFDM parameter $\lambda_{\text{DM}}$. This can be explained from the spacetime metric in expression (6) for Euler-Heisenberg black hole surrounded by PFDM. The dark matter effects contribute to the spacetime metric function $f(r)$ via the $-\lambda_{\text{DM}}/r$ term (if we naively omit the slowly varying logarithmic term $\ln\frac{r}{|\lambda_{\text{DM}}|}$ for simplicity), which is similar to the black hole’s mass term $2M/r$. This indicates that the PFDM parameter could play a role of “effective mass” in the spacetime metric  [^1] Strictly specking, because of the presence of logarithmic term, this “effective mass” should be $m_{\text{eff}}(r)=\lambda_{\text{DM}}\ln\frac{r}{|\lambda_{\text{DM}}|}$, rather than $m_{\text{eff}}=\lambda_{\text{DM}}$.. A positive PFDM parameter ($\lambda_{\text{DM}}>0$) contributes to a negative effective mass in the spacetime metric, which explains the rapid drop of the gravitational deflection angle when PFDM parameter becomes larger. Furthermore, combining the four panels of figure 3, it is also interesting to find that the gravitational deflection angles of light for Euler-Heisenberg black hole surrounded by PFDM (with nonzero electric charge $Q\neq 0$ and dark matter parameter $\lambda_{\text{DM}}\neq 0$) are smaller than those in Schwarzschild black hole cases (which correspond to the black solid line in the lower left panel of figure 3).
@@ 4.2 Time Delay of Light
The time delay results for Euler-Heisenberg black hole surrounded by PFDM are presented in figure 4. This figure plots the time delay of light measured by the conventional static time coordinate $t$ in spacetime metric (8). They are obtained numerically from expression (19) for light source and observer located at a finite distance region. In this figure, the time delays of light varied with respect to light source position $r_{\text{S}}$ are plotted. A similar plot can be carried out for time delays varied with respect to observer position $r_{\text{O}}$, while these results exhibit very similar behaviors with those in figure 4. So the plot of time delays varied with respect to observer position is not given in this work. In figure 4, the upper left and upper right panels highlight the influences from black hole electric charge $Q$ and the nonlinear electrodynamics effects (or QED effects) parameterized by $a$, while the lower panels highlight the influences form dark matter parameterized by $\lambda_{\text{DM}}$. In all four panels, the time delays of light all grow with the increasing of light source radius $r_{\text{S}}$, confirming that the time delay of light $\Delta T$ calculated using expression (19) is indeed a monotonically function of light source radius $r_{\text{S}}$ (or observer radius $r_{\text{O}}$).
Figure 4: The gravitational time delay of light for Euler-Heisenberg black hole surrounded by PFDM. The time delays of light (which are measured by the conventional static time coordinate $t$) are calculated numerically from expression (19) for finite distance light source and observer. This figure plots the time delay results, highlighting the influences coming from various black hole parameters. (a) The upper left panel shows the influences from black hole electric charge $Q$, with the nonlinear electrodynamics parameter $a=0.5M^{2}$ and dark matter parameter $\lambda_{\text{DM}}=0.2M$. (b) The upper right panel exhibits the influences from nonlinear electrodynamics effects / QED effects, with the black hole charge $Q=0.5M$ and PFDM parameter $\lambda_{\text{DM}}=0.2M$. (c) The lower left panel highlights the influences form dark matter in the absence of black hole charge and nonlinear electrodynamics effects (with $Q=0$ and $a=0$). (d) The lower right panel highlights the influences form dark matter in the presence of nonlinear electrodynamics effects, with the black hole charge $Q=0.5M$ and nonlinear electrodynamics parameter $a=0.5M^{2}$. In all four panels of this figure, the horizontal axis labels the position of light source varied from $r_{\text{S}}=100M$ to $r_{\text{S}}=1000M$, and the vertical axis labels the time delay of light $\Delta T$ in unit of black hole mass. The position of observer is chosen to be $r_{\text{O}}=200M$, and the impact parameter in photon orbit satisfies $b=10M$.
@@
Comparing the time delay results in figure 4 and the gravitational deflection results in figure 3, it is easy to observe that the black hole parameters (including the black hole electric charge $Q$, nonlinear electrodynamics parameter $a$, PFDM parameter $\lambda_{\text{DM}}$) influence time delays of light in a similar trend as those for gravitational deflection angles. Firstly, Euler-Heisenberg black hole with a larger electric charge results in a reduction of time delay, as shown in the upper left panel of figure 4. Secondly, the nonlinear electrodynamics effects (parameterized by $a$) do not have obvious influences on time delay results, as indicated in the upper right panel of figure 4. Thirdly, the most significant impacts on time delay coming from the PFDM, which are much greater than those from electric charge and nonlinear electrodynamics. From the lower panels of figure 4, the time delay of light changes rapidly as the PFDM parameter varies. Particularly, the time delay of light could undergo a rapid reduction for larger PFDM parameters, compared with those in the absence of dark matter (with $\lambda_{\text{DM}}=0$). This can be explained by the “effective mass” role of PFDM played in the spacetime metric, the same as we have elaborated in subsection 4.1 for gravitational deflection angle cases. A positive PFDM parameter $\lambda_{\text{DM}}>0$ contributes to a negative effective mass in the spacetime metric, which explains the rapid reduction of time delay for larger PFDM parameter values. Furthermore, combining the four panels of figure 4, it is worth noting that the time delays of light for Euler-Heisenberg black hole surrounded by PFDM (with nonzero electric charge $Q\neq 0$ and dark matter parameter $\lambda_{\text{DM}}\neq 0$) are smaller than those in Schwarzschild black hole cases (which correspond to the black solid line in the lower left panel of figure 4).
4.3 Massive Object’s Bound Orbits and Precession Angle
@@
Figure 5: The precession angles and trajectories of massive particle’s bound orbits moving around the Euler-Heisenberg black hole in the presence of PFDM. The precession angles $\Delta\phi_{\text{precession}}$ plotted in this figure are calculated numerically from expression (24) for a given conserved angular momentum value, and the trajectories of orbits are obtained by numerically solving the differential equation (23). The left panels shows the variation of precession angle as the changes of total energy parameter $\varepsilon_{\text{energy}}=\frac{E^{2}-1}{2}$, while the right panels illustrate the trajectories of particle orbits (It should be noted that we have substituted $1/2$ in the both side of equation (10) such that the energy parameter vanished at infinity, namely $\lim_{r\to\infty}\varepsilon_{\text{energy}}=0$). (a) The upper left panel plots the precession angle of orbits affects by black hole charge $Q$. (b) The upper right panel illustrates the trajectories of orbits for different black hole charge values when total energy equals to $\varepsilon_{\text{energy}}=-0.01$ (c) The middle left panel shows the precession angle of orbits affects by dark mater (parameterized by $\lambda_{\text{DM}}$). (d) The middle right panel illustrates the trajectories of orbits for different PFDM parameter values with total energy to be $\varepsilon_{\text{energy}}=-0.01$ (however, in the case of $\lambda_{\text{DM}}=0.2M$, the bound orbit does not exist because the local minimum of effective potential $V_{\text{min}}<\varepsilon_{\text{energy}}=-0.01$, see table 1) (e) The lower left panel shows the precession angle of orbits influenced by nonlinear electrodynamics effects / QED effects (parameterized by $a$). (f) The lower right panel illustrates the trajectories of orbits for different nonlinear electrodynamics parameter values with total energy to be $\varepsilon_{\text{energy}}=-0.01$
@@
The figure 5 displays numerical results on massive object’s bound orbits around the Euler-Heisenberg black hole surround by PFDM, including the precession angles in the left panels and the trajectories of orbits in the right panels. The trajectories of massive object’s orbits are obtained by numerically solving the ordinary differential equation (23), and the precession angles of orbits are calculated from the integral in expression (24). The upper, middle and lower panels of figure 5 highlight the influences from black hole electric charge, dark matter (parameterized by PFDM parameter $\lambda_{\text{DM}}$), and nonlinear electrodynamics effects / QED effects (parameterized by nonlinear electrodynamics parameter $a$) correspondingly. In this figure, we present results on the trajectory of orbits and their precession angles for a selected angular momentum value $J=4M$. The variation trend of precession angles correspond to other angular momentum values is similar to that for $J=4M$ presented in figure 5  [^2] For other angular momentum values, the proper region for conserved energy where bound orbits of massive objects could exist is different from those in figure 5..
@@
The results in figure 5 show that both black hole electric charge and PFDM can intensely influence the precession angle of massive object’s orbits in the vicinity of Euler-Heisenberg black hole. The upper panels of figure 5 suggests that a larger black hole charge reduces the precession angle of orbits. However, unlike the precession angle, the shape of massive object’s bound orbits are not violently influenced by black hole electric charge. In the upper right panel of figure 5, the shape and eccentricity of massive object’s orbits calculated for different black hole charge values do not exhibit large differences, until the electric charge grows to a value closer the extreme black hole cases (see the $Q=0.8M$ case, which corresponds to the dotted line in the upper right panel). On the other hand, the middle panels indicate that the PFDM have great impacts not only on the precession angle of orbits, but also on the shape of massive object’s bound orbits. The middle left panel tells that a larger PFDM parameter can significantly reduce the precession angle of bound orbits. The middle right panel shows that the shape, eccentricity and semimajor-axis length of massive object’s bound orbits exhibit notable differences when PFDM parameter takes different values. Moreover, in the middle left panel, the proper region for energy parameter $\varepsilon_{\text{energy}}=\frac{E^{2}-1}{2}$ which allows the existence of bound orbits is also reduced with the increasing of PFDM parameter. This is caused by the local minimum position of effective potential for massive objects. For any massive object’s bound orbits with precession, the total energy $\varepsilon_{\text{energy}}$ in this orbit must exceed the local minimum of effective potential. The local minimum values of effective potential for massive objects moving around Euler-Heisenberg black hole in the presence of PFDM (for a selected angular momentum value $J=4M$ and several PFDM parameter values) are listed in table 1. Furthermore, unlike the black hole electric charge and PFDM, the nonlinear electrodynamics effects / QED effects (parameterized by $a$) have very tiny impacts on the trajectories of massive object’s orbits and their procession angles, which can be directly observed from the lower panels of figure 5. This trend is similar to the cases of gravitational deflection and time delay, where the nonlinear electrodynamics effects also produce very tiny influences on these gravitational lensing observables.